# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ 3.1 Analysis of MB-RS-QVI
The proof follows immediately from Theorem 1 and Theorem 7 with the choice of $c=6$. ∎
By taking the limit $\beta\rightarrow 0$ leads to PAC-bounds in the risk-neutral case, but the resulting Q-value learning PAC bound is worse by a factor of $\frac{1}{1-\gamma}$ gheshlaghi2013minimax and the policy learning PAC-bound is worse by a factor of $\frac{1}{(1-\gamma)^{3}}$ agarwal_model-based_2020 .
Since the derivations so far lead to a PAC-bound on policy learning that is worse by a factor of $\frac{1}{(1-\gamma)^{2}}$ compared to the PAC bound on Q-value learning, we now provide an alternative derivation that leads to a PAC-bound with the same dependence on horizon. The bound achieved based on the Weissman bound however comes with a worse dependence on $S$, but in cases with long effective horizons and small state-space, this bound might be tighter. Another deficiency of this alternative bound is that it is also less interpretable in the sense that it explodes for $\beta\rightarrow 0$.
Theorem 3.
There exists a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$, and any MDP $M$ with $S$ states and $A$ actions, if the learner is allowed to make
$T=c\frac{SA(S+\log(SA/\delta))}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}$ (15)
calls to the generative model, then $\mathbb{P}(\|V^{*}-V^{\pi_{k}}\|\leq\varepsilon)\geq 1-\delta$.
Proof.
By Lemma 2, we can pick $k$ large enough so that $\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|\leq\varepsilon/3$. By the union bound we have that the two terms $\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|$ and $\|\widehat{V}^{\pi^{*}}-V^{*}\|$ can simultaneously be made smaller than $\varepsilon/3$ by sampling each state-action pair $N=8\frac{S+\log(\frac{2SA}{\delta})}{\tau^{2}}$ times where $\tau=\frac{\varepsilon}{3}\bigg[\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\bigg]^{-1}$ such that if we sample
$T=73\bigg[S+\log\bigg(\frac{SA}{\delta}\bigg)\bigg]\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}$ (16)
times in total, then we can ensure that for all $s$ simultaneously it holds that $V^{\pi_{k}}(s)>V^{*}-\varepsilon$. ∎
@@
State-of-the-art techniques for proving optimal upper bounds in the risk-neutral case are critically exploiting linearity of the expectation operator and are thus not readily available in the risk-sensitive case due to the non-linearity of the entropic risk measure.
4 Sample Complexity Lower Bounds
In this section, we provide two sample complexity lower bounds. The first one, presented in Theorem 4, concerns the sample complexity of learning the optimal $Q$-value function $Q^{*}$, whereas the second one, in Theorem 5, is on learning an optimal policy $\pi^{*}$. The proofs are both postponed to Appendix H.
Theorem 4 (Lower bound for learning $Q^{*}$).
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a $Q$-function $Q^{\mathcal{U}}$ and any $\delta\in(0,\frac{1}{4})$ and $\varepsilon\in\big(0,\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|\frac{1}{1-\gamma}})\big)$, the following holds: if the total number $T$ of transitions satisfies
$T\leq\frac{(S-2)A\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|\frac{1}{1-\gamma}}-3)}{|\beta|^{2}}\log\bigg(\frac{(S-2)A}{c_{2}\delta}\bigg)\,$,
then there is some MDP $M$ with $S$ states and $A$ actions for which $\mathbb{P}(\|Q^{*}_{M}-Q_{T}^{\mathcal{U}}\|>\varepsilon)\geq\delta$.
Theorem 5 (Lower bound for learning $\pi^{*}$).
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a policy $Q^{\mathcal{U}}$ and any $\delta\in(0,\frac{1}{4})$ and $\varepsilon\in\big(0,\frac{1}{50}\frac{\gamma}{|\beta|}(1-e^{-|\beta|\frac{1}{1-\gamma}})\big)$, it holds that if the total number $T$ of transitions satisfies
$T\leq\frac{(S-2)(A-1)\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|\frac{1}{1-\gamma}}-3)}{|\beta|^{2}}\log\bigg(\frac{(S-2)}{c_{2}\delta}\bigg),$
then there is some MDP $M$ with $S$ states and $A$ actions for which $\mathbb{P}(\|V^{*}_{M}-V^{\pi_{T}^{\mathcal{U}}}\|>\varepsilon)\geq\delta$.
The lower bounds in Theorems 4-5 establish that an exponential dependence of the sample complexity on the effective horizon $\frac{1}{1-\gamma}$ is unavoidable; more precisely, they assert that a scaling of $e^{|\beta|\frac{1}{1-\gamma}}$ is unavoidable in both $Q$-learning and policy-learning. Recalling the sample complexity bounds of MB-RS-QVI (Theorems 1-2), we observe a similar exponential dependence. However, there remains a gap of order $\frac{1}{(1-\gamma)^{2}}e^{|\beta|\frac{1}{1-\gamma}}$. This remains as an interesting open question as to whether closing this gap can be done via a more elegant analysis of MB-RS-QVI or it calls for more novel algorithmic ideas, but in any case the lower bounds show that risk-sensitive agents face a fundamentally harder problem than risk-neutral agents where the sample complexity is polynomial in $\frac{1}{1-\gamma}$. The proofs of lower bounds are given in Section H in the appendix, but we briefly sketch the ideas below for the case where the algorithm outputs a $Q$-function; the other case is proven using quite similar ideas.
We consider a class of hard-to-learn MDPs. In the class there are two absorbing states $s^{G}$ and $s^{B}$ where in $s^{G}$ the agent always recieves a reward of $R=1$ and in $s^{B}$ the agent always receives a reward of $R=0$ irrespective of the action taken. For all other state-action pairs $z$ the reward is zero and the only possible transition is to either of the states $s^{G}$ and $s^{B}$. $P(s^{G}|z)=q$ and $P(s^{B}|z)=1-q$, for some $q>0$. This construction critically allows us to calculate explicitly $Q^{*}(z)$ for a given parameter $q$ and for two different MDPs $M_{0},M_{1}$ in the class where $q_{0}=p$ and $q_{1}=p+\alpha$ for appropriately chosen values of $p$ and $\alpha$ we are able to ensure that $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$ which means that any specific algorithmic output $Q^{\mathcal{U}}(z)$ cannot be $\varepsilon$-close to both $Q^{*}_{M_{1}}(z)$ and $Q^{*}_{M_{0}}(z)$. We then show by a likelihood ratio argument that any algorithm $\mathcal{U}$ that is $(\varepsilon,\delta)$-correct on $M_{0}$, i.e. that $\mathbb{P}_{0}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$, will also satisfy that $\mathbb{P}_{1}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$ provided that the algorithm does not try out $z$ enough times on $M_{0}$ and exactly because $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$, the event $\{|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon\}$ is disjoint from the event on being $\varepsilon$-close to $Q_{M_{1}}^{*}$. The final part of the proof is to exploit that the different state-action pairs contain no information about each other which allows for an independence argument for the estimation of $Q^{\mathcal{U}}(z)$ and $Q^{\mathcal{U}}(z^{\prime})$ for $z\neq z^{\prime}$. While doing this analysis, we fix an inaccuracy in the proof of Lemma 17 in gheshlaghi2013minimax that arises where they lower-bound the likelihood ratio of two Bernoulli random variables with biases $p\geq\frac{1}{2}$ and $p+\alpha$ on a high probability event. We also mention that we extend the result to hold for $p<\frac{1}{2}$.
For policy learning we consider almost the same class of MDPs but augment with a known state $a_{0}$ that is used in the analysis.
@@ 4 Sample Complexity Lower Bounds
It is worth remarking that the best lower bound in the risk-neutral setting is derived in gheshlaghi2013minimax using a richer construction than above. However, with a risk-sensitive learning objective, the optimal state-action value function in the construction of gheshlaghi2013minimax does not admit an analytical solution, which is needed for the delicate tuning of the transition probabilities.
5 Conclusion and Future Works
We have studied the sample-complexity of learning the optimal $Q$-function and that of learning an optimal policy in finite discounted MDPs, where the agent has recursive risk-preferences given by the entropic risk measure and has access to a generative model. We introduced an algorithm, called MB-RS-QVI, and derived PAC-type bounds on its sample complexity for both learning which have derived bounds from analyzing the MB-RS-QVI algorithm that uses the model given by the plug-in estimator from samples generated by a simulator. We also derive lower bounds. The upper bounds show that PAC-learning is possible but the lower bounds show that dependence on $e^{|\beta|\frac{1}{1-\gamma}}$ is unavoidable and thus that learning is fundamentally harder for risk-senstive agents relative to risk-neutral agents. The bounds that we derive on the sample complexity of learning the optimal $Q$-value are of order
$\mathcal{O}\bigg(\log(SA/\delta)\big)\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\bigg),\qquad\Omega\bigg(\frac{(S-2)A}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log((S-2)A/\delta)\bigg)$ (17)
while the bounds we derive on the sample complexity of learning an optimal policy are of order
$\mathcal{O}\bigg(\log(SA/\delta)\big)\frac{SA}{\varepsilon^{2}(1-\gamma)^{4}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\bigg),\qquad\Omega\bigg(\frac{(S-2)(A-1)}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log((S-2)/\delta)\bigg)$ (18)
where we also give the alternative bound of order
$\qquad\mathcal{O}\bigg(\big(S+\log(SA/\delta)\big)\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\bigg)$ (19)
which might be tighter in cases of long horizon and small state space. These constitute the first bounds, to our knowledge, on the sample complexities of entropic risk-sensitive agents in the discounted MDP setting. The upper and lower bounds derived in this paper leave open gaps in $\frac{1}{1-\gamma}$. Since the constructions in the lower bounds are not the ones used in the tightest lower bounds derived in the risk-neutral setting, one possibility is that the lower bounds can be improved by considering a more carefully chosen set of hard-to-learn MDPs with the challenge being to control the gap in $V$-values or $Q$-values under different parameters. Also since the plugin-estimator model-based QVI algorithm is provably optimal in the risk-neutral setting, we believe that this might also be the case for risk-sensitive agents but that more tools are needed to develop a more careful analysis. Another future direction is that of developing model-free algorithms for this setting and analyzing their statistical efficiency. Another interesting research direction is to consider function approximation, as in zhou2021provably . As other future directions, one may consider more complicated RL settings such as offline RL rashidinejad2021bridging , where data is collected under a fixed (but unknown) behavior policy, and online RL strehl_analysis_2008 ; lattimore_near-optimal_2014 , where the agent’s learned policy impacts the data collection process. And finally one may also consider the problem where the learner have non-recursive risk-preferences instead.

### v3

@@ 6.1.1 Proof of Theorem 2
Let $\varepsilon>0$ and $\beta\neq 0$. In view of the error decomposition in (9), to establish $\varepsilon$-policy-correctness it suffices to require: (i) $\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|\leq\varepsilon/3$, (ii) $\|\widehat{V}^{\pi^{*}}-V^{*}\|\leq\varepsilon/3$, and (iii) $\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|\leq\varepsilon/3$. For (i), we can have $\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|\leq\varepsilon/3$ using enough iterations of the optimization oracle as a consequence of Lemma 2. To control (ii) and (iii), let us define the event
$E:=\cap_{\pi\in\Pi}\big\{\|V^{\pi}-\widehat{V}^{\pi}\|\leq\varepsilon/3\}.$
We can show that $\mathbb{P}(E)\geq 1-\delta$ for sufficiently large $N$. Indeed, for a given $\pi\in\Pi$, an application of Lemma 3 and Lemma 6.1, identical to the treatment in the proof of Theorem 2, shows that $\|V^{\pi}-\widehat{V}^{\pi}\|\leq\varepsilon/3$ with probability at least $1-\delta/|\Pi|$ if $N\geq\frac{1}{2\tau^{2}}\big(1-e^{-\beta/(1-\gamma)}\big)^{2}\log(2SA|\Pi|/\delta)$ with $\tau=\frac{\varepsilon(1-\gamma)|\beta|}{3\gamma}e^{-|\beta|/(1-\gamma)}$. Hence,
$\mathbb{P}(E^{c})=\mathbb{P}\big(\exists\pi\in\Pi:\|V^{\pi}-\widehat{V}^{\pi}\|>\varepsilon/3\big)\leq\sum_{\pi\in\Pi}\mathbb{P}\big(\|V^{\pi}-\widehat{V}^{\pi}\|>\varepsilon/3\big)\leq\delta,$
so that $\mathbb{P}(E)\geq 1-\delta$. It is evident that conditioned on $E$, (ii) and (iii) holds. Hence, with probability greater than $1-\delta$, $\varepsilon$-policy-correctness is maintained using a total number of model calls of
$T_{1,\delta}=\frac{9SA\gamma^{2}}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|/(1-\gamma)}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{2SA|\Pi|}{\delta}\bigg)\,.$ (11)
To establish the second bound, first observe that Theorem 1 implies that $\|\widehat{V}^{\pi_{k}}-V^{*}\|\leq\varepsilon$ with probability exceeding $1-\delta$ if $T\geq\frac{2SA\gamma^{2}}{\varepsilon^{2}(1-\gamma)^{2}|\beta|^{2}}\big(e^{|\beta|/(1-\gamma)}-1\big)^{2}\log\big(\frac{2SA}{\delta}\big)$; this is verified by noting that
$\|\widehat{V}^{\pi_{k}}-V^{*}\|=\max_{s}\big|\max_{a}\widehat{Q}^{\pi_{k}}(s,a)-\max_{a}Q^{*}(s,a)\big|\leq\max_{s,a}\big|\widehat{Q}^{\pi_{k}}(s,a)-Q^{*}(s,a)\big|=\|Q_{k}-Q^{*}\|.$
To proceed, we make use of the following lemma, which is proven in Appendix D:
{restatable}
lemmaDeteriorationGreedyPolicy Let $\alpha>0$. Let $\overline{V}\in\mathbb{R}^{S}$ be a value function obeying $\|V^{*}-\overline{V}\|<\alpha$, and $\pi^{G}:=\arg\!\max_{a}[R(s,a)+\gamma\rho_{s,a}(\overline{V}(s^{\prime}))]$ be a greedy policy with respect to $\overline{V}$. Then, $\|V^{*}-V^{\pi_{G}}\|\leq\frac{2\gamma}{1-\gamma}\alpha$ .
Choosing $\overline{V}=\widehat{V}^{\pi_{k}}$, we have shown that $\|\widehat{V}^{\pi_{k}}-V^{*}\|\leq\varepsilon$ with probability $1-\delta$. Further, note that $\pi_{k}$ by construction is the greedy policy with respect to $\widehat{V}^{\pi_{k}}$. Therefore, by Lemma 6.1.1, the true value of $\pi_{k}$ satisfies, with probability at least $1-\delta$,
$\|V^{*}-V^{\pi_{k}}\|\leq\frac{2\gamma}{1-\gamma}\varepsilon,$
which yields the following bound:
$T_{2,\delta}=\frac{8SA\gamma^{2}}{\varepsilon^{2}(1-\gamma)^{4}}\bigg(\frac{e^{|\beta|/(1-\gamma)}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{2SA}{\delta}\bigg)\,.$ (12)
Final Bound.
To derive the final bound, we put together (11) and (12), while suitably adjusting the error probabilities. Therefore, $\varepsilon$-policy-correctness is guaranteed with probability exceeding $1-\delta$ if $T\geq\min\{T_{1,\delta/2},T_{2,\delta/2}\}$. ∎
7 Sample Complexity Lower Bounds
In this section, we provide two sample complexity lower bounds. The first (Theorem 3) concerns value learning, whereas the second (Theorem 4) addresses policy learning.
Theorem 3 (Lower bound for value learning).
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a $Q$-value $Q^{\mathcal{U}}$, any $\delta\in(0,\frac{1}{4})$, and $\varepsilon\in\big(0,\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|/(1-\gamma)})\big)$, the following holds: if the total number $T$ of samples satisfies
$T\leq\frac{SA\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|/(1-\gamma)}-3)}{|\beta|^{2}}\log\bigg(\frac{SA}{c_{2}\delta}\bigg)\,$,
then there exists some MDP $M$ with $S$ states and $A$ actions for which $\mathbb{P}(\|Q^{*}_{M}-Q_{T}^{\mathcal{U}}\|>\varepsilon)\geq\delta$.
Theorem 4 (Lower bound for policy learning).
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a policy $\pi^{\mathcal{U}}_{T}$, any $\delta\in(0,\frac{1}{4})$, and $\varepsilon\in\big(0,\frac{1}{50}\frac{\gamma}{|\beta|}(1-e^{-|\beta|/(1-\gamma)})\big)$, it holds that if the total number $T$ of samples satisfies
$T\leq\frac{SA\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|/(1-\gamma)}-3)}{|\beta|^{2}}\log\bigg(\frac{S}{c_{2}\delta}\bigg),$
then there exists some MDP $M$ with $S$ states and $A$ actions for which $\mathbb{P}(\|V^{*}_{M}-V^{\pi_{T}^{\mathcal{U}}}\|>\varepsilon)\geq\delta$.
While analogous policy learning lower bounds are often stated in the risk-neutral literature, explicit proofs are typically omitted, to the best of our knowledge. For completeness, we provide a detailed, step-by-step derivation, emphasizing its close connection to the corresponding value-learning lower bound as well as the subtle differences that arise in the final guarantee.
Theorems 34 establish that an exponential dependence on the effective horizon $1/(1-\gamma)$ in the sample complexity is unavoidable under both value and policy learning. These bounds cover both risk-averse ($\beta>0$) and risk-seeking ($\beta<0$) agents, providing strong impossibility results for recursive ERM. Comparing with the sample complexity bounds of MB-RS-QVI (Theorems 12), we observe a similar exponential dependence; however, a gap of order $\frac{1}{(1-\gamma)^{2}}e^{|\beta|/(1-\gamma)}$ remains. Closing this gap may require either a refined analysis of MB-RS-QVI or more sophisticated algorithmic ideas. Nevertheless, these lower bounds confirm that risk-sensitive RL under ERM is fundamentally harder in the worst case than the risk-neutral setting, where minimax sample complexity scales polynomially with $1/(1-\gamma)$.
$z_{i}$$s^{G}$$s^{B}$$q_{i}$$1-q_{i}$R=1R=1R=0R=0
Figure 1: Hard-to-learn MDP construction
Proof sketch.
The proofs are provided in Appendix F; here we sketch the main ideas for value learning (Theorem 3). The construction involves a class of hard-to-learn MDPs (Figure 1) with two absorbing states $s^{G}$ and $s^{B}$, yielding rewards $R=1$ and $R=0$ under any action, respectively. All other state-action pairs $z$ give zero reward and transition only to $s^{G}$ or $s^{B}$ with probability $P(s^{G}|z)=q$ and $P(s^{B}|z)=1-q$, for some $q>0$. This construction critically allows us to calculate explicitly $Q^{*}(z)$ for a given parameter $q$ and for two different MDPs $M_{0},M_{1}$ in the class, where $q_{0}=p$ and $q_{1}=p+\alpha$ for appropriately chosen values of $p$ and $\alpha$. It is key to choose them in a way to ensure that $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$, which means that any specific algorithmic output $Q^{\mathcal{U}}(z)$ cannot be $\varepsilon$-close to both $Q^{*}_{M_{1}}(z)$ and $Q^{*}_{M_{0}}(z)$. We then show by a likelihood ratio argument that any algorithm $\mathcal{U}$ that is $(\varepsilon,\delta)$-correct on $M_{0}$, i.e. that $\mathbb{P}_{0}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$, will also satisfy that $\mathbb{P}_{1}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$ provided that the algorithm does not try out $z$ enough times on $M_{0}$ and exactly because $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$, the event $\{|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon\}$ is disjoint from the event on being $\varepsilon$-close to $Q_{M_{1}}^{*}$. The final part of the proof is to exploit that the different state-action pairs contain no information about each other, which allows for an independence argument for the estimation of $Q^{\mathcal{U}}(z)$ and $Q^{\mathcal{U}}(z^{\prime})$ for $z\neq z^{\prime}$.
We note that, in the course of this analysis, we also correct a minor issue in Lemma 17 of [30]. Specifically, the issue arises in the derivation of a lower bound on the likelihood ratio between two Bernoulli distributions with means $p\geq\frac{1}{2}$ and $p+\alpha$ on a high probability event, for $p>\frac{1}{2}$. Additionally, we establish a corresponding bound for $p<\frac{1}{2}$, which may be of independent interest. For policy learning, a similar construction is used, augmented with a known action $a_{0}$ to facilitate the analysis.
Algorithmic intuition.
The above construction also provides insight into how a model-based algorithm such as MB-RS-QVI behaves on these instances. Since all rewards are zero except in the absorbing states, learning reduces to estimating the transition probability $q=P(s^{G}|z)$ for each state-action pair $z$. The algorithm therefore forms an empirical estimate $\hat{q}$ and computes $Q$-values based on this estimate. On the hard instances $M_{0}$ and $M_{1}$, the true probabilities differ only by a small amount ($q_{0}=p$ vs. $q_{1}=p+\alpha$). When the number of samples is limited, the empirical estimate $\widehat{q}$ will typically not be accurate enough to reliably distinguish between these two cases. As a result, the algorithm may construct a model that is statistically consistent with both $M_{0}$ and $M_{1}$, leading to $Q$-value estimates that are necessarily inaccurate for at least one of them.
@@ 7 Sample Complexity Lower Bounds
The best lower bound in the risk-neutral setting is derived in [30] using a richer construction than above. However, with a risk-sensitive learning objective, the optimal Q-value function in that construction does not admit an analytical solution, which is crucial for tuning transition probabilities and deriving our bounds.
@@
We note that the bound becomes vacuous for $|\beta|\leq(1-\gamma)\log(3)$. This is partly due to a final approximation introduced to make the bound more interpretable; importantly, this approximation does not affect the dependence on $|\beta|$ for large $|\beta|$. Even without this approximation, the bound still becomes vacuous as $\beta\rightarrow 0$. This behavior arises because, in this limit, $p\to 1$ or $p\to 0$ depending on the direction of the limit. Since our information-theoretic argument yields a sample complexity proportional to p(1p)p(1-p), the bound vanishes in this regime.
8 Numerical experiments
We conduct a numerical experiments to showcase the empirical performance of MB-RS-QVI. We consider a RiverSwim MDP [70] (Figure 22(a)) with $8$ states with discount factor $\gamma=0.95$. In this MDP, there two actions in each state, corresponding to moving ‘left’ or ‘right’. The rewards are zero, except in two places: a reward of $0.05$ (low reward) under ‘left’ in the left-most state ($s=1$), and reward of $1$ (high reward) under ‘right’ in the right-most state ($s=L$). The low reward is easy to access because of the actions with deterministic transitions. The agent has a risk-sensitive objective defined using recursive ERM. To showcase the performance of MB-RS-QVI, we consider three different values of $\beta\in\{0,1,1.25\}$; we recall that $\beta=0$ corresponds to the risk-neutral case. For each value of $\beta$, we generated datasets of increasing sizes $T\in\{160,320,\ldots,1600\}$ (multiples of $SA=16$) by sampling each state-action pair in the MDP. The MB-RS-QVI algorithm was run on each dataset to produce an output policy $\hat{\pi}$. This procedure was run for $1000$ independent runs.
Figure 2 depicts $\|V^{*}-V^{\hat{\pi}}\|$ averaged over the $1000$ runs for the three values of $\beta$. The true value of $\hat{\pi}$ is computed using a value iteration procedure based on (7). In this figure, one may consider a particular value of $\varepsilon$ to observe the number of samples needs to output an $\varepsilon$-optimal policy. It is evident that this number of samples increases as $\beta$ increases. Furthermore, it is evident that even a small increase in $\beta$ (from $1$ to $1.25$) leads to a large increase in the number of samples to learn an $\varepsilon$-optimal policy. We also note that when the output policy $\widehat{\pi}$ matches the optimal policy for all the runs, the curve would hit the horizontal axis. These results showcases that learning a near-optimal policy under ERM could be substantially more demanding in terms of data even in such a rather simple MDP, and the data efficiency may be severely impacted as the risk parameter $\beta$ increases.
Figure 2: (a) The $L$-state RiverSwim MDP [70]; (b) Policy learning error $\|V^{*}-V^{\hat{\pi}}\|$ under MB-RS-QVI.
9 Concluding Remarks
We studied the sample complexities of value learning and policy learning in finite discounted MDPs, where the agent exhibits recursive risk preferences modeled via the ERM and has access to a generative model. The generative model setting is commonly used in theoretical RL as it provides a clean framework in which the statistical difficulty of the problem can be isolated and precisely characterized. In particular, it removes effects due to exploration and data collection, allowing us to focus on the intrinsic difficulty induced by the risk-sensitive objective. We introduced a model-based algorithm, MB-RS-QVI, and derived PAC-type bounds on its sample complexity. In addition, we established sample complexity lower bounds for both policy and value learning. These lower bounds reveal that an exponential dependence on the horizon $1/(1-\gamma)$ is unavoidable in the worst case, demonstrating that this setting is fundamentally harder than the risk-neutral case. To our knowledge, these constitute the first upper and lower bounds for this setting. The bounds are tight in $S$, AA, $\delta$, and $\varepsilon$, but gaps remain in the dependence on $1/(1-\gamma)$.
Closing these remaining gaps leads to interesting directions for future work. Firstly, our lower bound is valid for $|\beta|>(1-\gamma)\log(3)$, and it would therefore be interesting to derive a lower bound that would uncover the risk-neutral lower bound as $\beta$ tends to zero. It is not immediately clear whether this requires improving the upper bound, the lower bound, or both. Designing hard MDP instances for lower bounds is subtle: they should be difficult to learn, yet still tractable enough to compute value functions; this is a challenge amplified in the risk-sensitive setting. The plug-in estimator used in MB-RS-QVI is provably optimal in the risk-neutral case, raising the natural question of whether this optimality extends to ERM. Another promising direction is the development of model-free algorithms for this setting and the analysis of their statistical efficiency.
It would also be valuable to extend the study to MDPs with function approximation (e.g., [78]), as well as more complex RL settings, including offline RL [61], where data is collected under a fixed (but unknown) behavior policy, and online RL [70, 45], where the agent’s learned policy directly impacts the data collection process. A representative next step is offline RL, where, under a model-based approach, plug-in estimators combined with pessimism-style corrections are often used (e.g., [48]). Some technical tools developed in this paper may be of independent interest and could potentially be useful in analyzing related offline RL settings.